OpenStudy (anonymous):

Solve by the elimination method 0.3x-0.2y=4 0.4x+0.5y=1 an ordered pair

OpenStudy (anonymous):

x=220/23 and y =- 130/23 do u want me to show u how to get it?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

multiply the first equation by 5 and the second equation by 2

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

i just put your answer in, and it was incorrect

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

my mistake

OpenStudy (anonymous):

:) no prob